# GMValuator: Similarity-based Data Valuation for Generative Models

### v7

@@ Appendix C Different Generated Data Sizes
Since our value function $\phi_{i}$ (Eq. (3)) for training data $x_{i}$ is computed by averaging over generated samples, it is expected that the sensitivity of $\phi_{i}$ is connected to the size $m$ of the investigated generated sample. To explore the influence of generated data size $m$ on the utilization of GMValuator, we perform sensitivity testing on MNIST, CIFAR10 using generative models GAN [21], and Diffusion models [39] as depicted below. The dataset, model and used $k$ are denoted under each subfigure of Figure 7. First, we generate a varying number of samples from the same class. Specifically, we consider four different sample sizes, denoted by $m$, which are given by 1, 10, 30, 50. Next, we evaluate the GMValuator using parameter C1, and this evaluation is performed for each of the aforementioned values of $m$. Subsequently, we conduct the experiment 10 times using GMValuator (No-Rerank), each time with different $m$-sized generated data samples from the same class. The results are presented as the mean and standard deviation ($\rho$) for accuracy, taken over these 10 runs. The results shown in Figure 7 imply that varying $m$ does not yield notable differences in mean accuracy and increasing the number of generated samples $m$ leads to more stable and consistent results.
Appendix D Alternative Distance Metric
@@
We suggest utilizing Learned Perceptual Image Patch Similarity (LPIPS) [28] or DreamSim [27] as the distance metric $d$ during the re-ranking phase. This enables understanding the perceptual dissimilarity between generated and real data points. These metrics are applied in the image embedding space derived from pretrained models. Alternatively, distance measurement can be based on pixel space, such as employing $l2$-distance. Notably, we find that distances calculated in the input pixel space yield comparable outcomes as shown in Table VI. This implies that our method GMValuator could be flexible to the difference choices of distance metrics and the selection can depends on data prior.
Appendix E Additional Experimental Details
E-A Datasets
@@
We conduct the generation tasks in the experiments on benchmark datasets (i.e., MNIST [34] and CIFAR [35]), face recognition dataset (i.e., CelebA [36]), and high-resolution image dataset AFHQ [37] and FFHQ [38].
@@
E-B Architecture of Generative models
In our experiments, we leverage different generative models in the class of GAN, VAE and diffusion model. We utilize $\beta$-VAE for both MNIST and CIFAR-10 datasets while a simple GAN is conducted on MNIST. BigGAN and $\beta$-VAE are also conducted on CIFAR-10. We list the architecture details for these generative models in Table VII. StyleGAN is used for high-resolution dataset AFHQ and FFHQ. CelebA uses Diffusion-StyleGAN [33], for which we use the exact architecture in their open sourced code.
Appendix F Discussion on the Possible Applications
TABLE VII: The architecture of $\beta$-VAE and BigGAN.
GAN for MNIST

### v8

@@ Appendix F Different Generated Data Sizes
Since our value function $\phi_{i}$ (Eq. equation 5) for training data $x_{i}$ is computed by averaging over generated samples, it is expected that the sensitivity of $\phi_{i}$ is connected to the size $m$ of the investigated generated sample. To explore the influence of generated data size $m$ on the utilization of GMValuator, we perform sensitivity testing on MNIST, CIFAR10 using generative models GAN Goodfellow et al. (2020), and Diffusion models Dhariwal & Nichol (2021) as depicted below. The dataset, model and used $k$ are denoted under each subfigure of Figure 7. First, we generate a varying number of samples from the same class. Specifically, we consider four different sample sizes, denoted by $m$, which are given by 1, 10, 30, and 50. Next, we evaluate the GMValuator using parameter C1, and this evaluation is performed for each of the aforementioned values of $m$. Subsequently, we conduct the experiment 10 times using GMValuator (No-Rerank), each time with different $m$-sized generated data samples from the same class. The results are presented as the mean and standard deviation ($\rho$) for accuracy, taken over these 10 runs. The results shown in Figure 7 imply that varying $m$ does not yield notable differences in mean accuracy and increasing the number of generated samples $m$ leads to more stable and consistent results.
Appendix G Alternative Embedding Approaches
In the step of efficient similarity matching, the embedding $f_{e}$ is exclusively used in the recall phase for retaining $n$ samples with non-minimal contributions. Apart from using CLIP for embedding, we also investigate the influence for the results when using other embedding methods. We present results in Table 9 using more embedding methods, showing similar performance with CLIP with “Rerank”. The reason for this results is that we implement re-ranking that relies on the image space rather than the embedding space, which allows us to derive accurate top $k$ contributors from $n$ samples $n\gg k$. Thus, the embedding could tolerate some noise in the recall phase.
Table 9: Comparison of different embedding methods.
MNIST (%) No-Rerank $l_{2}$-distance LPIPS DreamSim
CLIP 86.41 87.76 88.78 88.78
Alexnet 79.77 84.82 86.51 88.72
Densenet 80.47 86.77 87.97 89.35
Appendix H Alternative Distance Metric
@@
We suggest utilizing Learned Perceptual Image Patch Similarity (LPIPS) Zhang et al. (2018) or DreamSim Fu et al. (2023) as the distance metric $d$ during the re-ranking phase. This enables to understand the perceptual dissimilarity between generated and real data points. These metrics are applied in the image embedding space derived from pre-trained models. Alternatively, distance measurement can be based on pixel space, such as employing $l2$-distance. Notably, we find that distances calculated in the input pixel space yield comparable outcomes as shown in Table 7. This implies that our method GMValuator could be flexible to the different choices of distance metrics and the selection can depend on data prior.
Appendix I Necessity for calibration
For challenges 2 and 3, we utilize image quality assessment and establish a non-zero scores rule for calibration in GMValuator. To better understand the impact and necessity of calibration, we compare the distribution of the top 1, 2, and 3 contributors’ scores with (w) and without (w/o) calibration conducted on MNIST in Figure 9. It is evident that scores without calibration are generally higher than with calibration. We also extend C3 and perform an ablation study in Figure 9 on scenarios where the generated samples include low-quality outputs. The results show that the OOD (out-of-distribution) training samples’ value rank by GMValuator without (w/o) calibration is smaller, indicating significant bias and poor performance.
Figure 8: Sore Distribution
Figure 9: Value Rank
Appendix J Additional Experimental Details
J.1 Justification of Experiment Setup
Table 10: Justification of Experiment Setup. The selection of datasets and models was based on three critical factors that guided the process. Firstly, attribute labels were required to evaluate C2 effectively. Secondly, benchmark datasets were meticulously chosen to ensure a fair comparison with baselines while also taking into account computational costs (C3 and C4). Finally, the selected generative models are powerful enough to generate good-quality data for the datasets.
Dataset Model Baseline Label Requirements
C1 MNIST, CIFAR-10 VAE VAE-TracIn Kong & Chaudhuri (2021) Class labels
Diffusion, GAN - Class labels
ImageNet Masked Diffusion Transformer Gao et al. (2023) - Class labels
C2 CelebA Diffusion-StyleGAN - Attribute labels
AFHQ, FFHQ StyleGAN - -
C3 MNIST DCGAN IF4GAN Terashita et al. (2021) Class labels
C4 MNIST, CIFAR-10 VAE VAE-TracIn Kong & Chaudhuri (2021) Class labels
MNIST DCGAN IF4GAN Terashita et al. (2021) Class labels
We provide a detailed justification in Table 10 for our experiment setup from C1 to C4.
  • C1. In Identical Class Test, the baseline method is VAE-TracIn Kong & Chaudhuri (2021), which can find the most influenced instances in the training dataset. Since VAE-TracIn is the model-specific method, we only need to compare it with GMValuator when VAE model is used. Besides, all the datasets used in C1 should have the class labels. Considering the computational demands detailed in VAE-TracIn Kong & Chaudhuri (2021), the runtime complexity of VAE-TracIn correlates with the number of network parameters and the size of the dataset. Therefore, our analysis primarily utilizes simpler benchmark datasets such as MNIST and CIFAR-10 for comparative evaluations with VAE-TracIn.
  • C2. In extending C1, we use image attributes to supplant the concept of class for datasets lacking class labels. For datasets with attribute labels, quantified experiments are feasible, as demonstrated in Table 3 and Table 8. For datasets without attribute labels, we employ visualized experiments.
  • C3. IF4GAN, a model-specific method for GAN, serves as the baseline for measuring data value. Adhering to the settings outlined in Terashita et al. (2021) and considering the impractical computational costs, we conduct experiments on the same dataset (MNIST) in Terashita et al. (2021) using DCGAN.
  • C4. We present a comparison of the efficiency of GMValuator against baseline methods. In accordance with the settings used for VAE-TracIN and IF4GAN in Kong & Chaudhuri (2021), we report the efficiency results for C1 on the MNIST and CIFAR-10 datasets, and for C3 on MNIST.
J.2 Datasets
@@ Appendix I Necessity for calibration
We conduct the generation tasks in the experiments on benchmark datasets (i.e., MNIST LeCun et al. (1998) and CIFAR Krizhevsky et al. (2009)), face recognition dataset (i.e., CelebA Liu et al. (2018)), high-resolution image dataset AFHQ Choi et al. (2020) and FFHQ Karras et al. (2019), large-scale image dataset ImageNet Deng et al. (2009).
@@
ImageNet. ImageNet is a large-scale image dataset, which contains over 14 million images and is categorized into more than 20,000 classes.
J.3 Architecture of Generative models
In our experiments, we leverage different generative models in the class of GAN, VAE and diffusion models. We utilize $\beta$-VAE for both MNIST and CIFAR-10 datasets while a simple GAN is conducted on MNIST. BigGAN and $\beta$-VAE are also conducted on CIFAR-10. We list the architecture details for these generative models from Table 11 to Table 13. StyleGAN is used for high-resolution datasets AFHQ and FFHQ. CelebA uses Diffusion-StyleGAN Wang et al. (2022), for which we use the exact architecture in their open-sourced code. In addition, Masked Diffusion Transformer, as introduced by Gao et al. Gao et al. (2023), is applied to the ImageNet.
Table 11: The architecture of GAN for MNIST.
@@
Appendix K Discussion on the Possible Applications
Appendix L Current Limitation and Future Directions
The limitation of this work is that it only measures data value for vision-related generative models and conducts experiments exclusively within the field of computer vision. However, this does not mean that GMValuator cannot be easily adapted to Natural Language Processing (NLP) fields, given its core idea of similarity matching. In the future, we should extend GMValuator to NLP and assess the data value for language-related generative models, such as large language models (LLMs).
Appendix M Omitted proofs
We follow Just et al. (2023) to prove the theorem. Firstly, we give several assumptions that will be used in later proof.
Assumption M.1
Following Assumption 2.3, given a distance function $d(\cdot,\cdot)$ between , we defined the coupling between $\mathcal{X}_{(T|f)}$ and $\mathcal{X}_{(S^{*}|f)}$ as $\pi^{*}$:
$\pi^{*}:=\operatorname*{arg\,inf}_{\pi\in\Pi(\mathcal{X}_{(T|f)},\mathcal{X}_{(S^{*}|f)})}\mathbb{E}_{(x_{T},x_{S^{*}})\sim\pi}d(x_{T},x_{S^{*}})$ (7)
It is easy to see that all joint distributions defined above are couplings between the corresponding distribution pairs. Then, following Just et al. (2023) we prove the main Theorem.
Theorem M.2
(Restated of Theorem 2.4.) Let $f_{S^{*}}^{{}^{\prime}}:\mu\rightarrow\mathcal{A}=\{0,1\}^{V}$ be the model trained on the optimal contributor dataset $S^{*}$. Following Assumption 2.3, if the contributors are corresponding to the given generated data $\hat{X}$, we have:
$\mathbb{E}_{x\sim\mu_{T}}\left[\mathcal{L}\left(f(x),f_{S^{*}}^{{}^{\prime}}(x)\right)\right]-\mathbb{E}_{x\sim\mu_{S^{*}}}\left[\mathcal{L}\left(f(x),f_{S^{*}}^{{}^{\prime}}(x)\right)\right]$ (8)
$\leq k\epsilon\cdot\!\left[d_{W}(\mathcal{X}_{(T|f)},\mathcal{X}_{(\hat{X}|f)})\!+\!d_{W}(\mathcal{X}_{(S^{*}|f)},\mathcal{X}_{(\hat{X}|{f})})\right]$
Proof M.3
$\mathbb{E}_{x\sim\mu_{T}}\left[\mathcal{L}\left(f(x),f_{S^{*}}^{{}^{\prime}}(x)\right)\right]$ $=\mathbb{E}_{x\sim\mu_{T}}\left[\mathcal{L}\left(f(x),f_{S^{*}}^{{}^{\prime}}(x)\right)\right]$
$-\mathbb{E}_{x\sim\mu_{S^{*}}}\left[\mathcal{L}\left(f(x),f_{S^{*}}^{{}^{\prime}}(x)\right)\right]+\mathbb{E}_{x\sim\mu_{S^{*}}}\left[\mathcal{L}\left(f(x),f_{S^{*}}^{{}^{\prime}}(x)\right)\right]$
$\leq\mathbb{E}_{x\sim\mu_{S^{*}}}\left[\mathcal{L}\left(f(x),f_{S^{*}}^{{}^{\prime}}(x)\right)\right]$
$+\left|\mathbb{E}_{x\sim\mu_{S^{*}}}\left[\mathcal{L}\left(f(x),f_{S^{*}}^{{}^{\prime}}(x)\right)\right]-\mathbb{E}_{x\sim\mu_{T}}\left[\mathcal{L}\left(f(x),f_{S^{*}}^{{}^{\prime}}(x)\right)\right]\right|$ (9)
We bound $\left|\mathbb{E}_{x\sim\mu_{S^{*}}}\left[\mathcal{L}\left(f(x),f_{S^{*}}^{{}^{\prime}}(x)\right)\right]-\mathbb{E}_{x\sim\mu_{T}}\left[\mathcal{L}\left(f(x),f_{S^{*}}^{{}^{\prime}}(x)\right)\right]\right|$ as follows:
$\left|\mathbb{E}_{x\sim\mu_{S^{*}}}\left[\mathcal{L}\left(f(x),f_{S^{*}}^{{}^{\prime}}(x)\right)\right]-\mathbb{E}_{x\sim\mu_{T}}\left[\mathcal{L}\left(f(x),f_{S^{*}}^{{}^{\prime}}(x)\right)\right]\right|$
$=\left|\int_{\mathcal{X}^{2}}\mathcal{L}\left(f(x_{S^{*}}),f_{S^{*}}^{{}^{\prime}}(x_{S})\right)-\mathcal{L}\left(f(x_{T}),f_{S^{*}}^{{}^{\prime}}(x_{T})\right)d\pi^{*}(x_{T},x_{S^{*}})\right|$
$=|\int_{\mathcal{X}^{2}}\mathcal{L}\left(f(x_{S^{*}}),f_{S^{*}}^{{}^{\prime}}(x_{S^{*}})\right)$
$-\mathcal{L}\left(f(x_{S^{*}}),f_{S^{*}}^{{}^{\prime}}(x_{T})\right)+\mathcal{L}\left(f(x_{S^{*}}),f_{S^{*}}^{{}^{\prime}}(x_{T})\right)-\mathcal{L}\left(f(x_{T}),f_{S^{*}}^{{}^{\prime}}(x_{T})\right)d\pi^{*}(x_{T},x_{S^{*}})|$
[…]
Then due to $k$-Lipschitzness of $\mathcal{L}$ and $\epsilon$-Lipschitzness of $f$, we can obtain:
[…]
$=k\epsilon d_{W}(\mathcal{X}_{(T|f)},\mathcal{X}_{(S^{*}|f)}),$
where the last step is due to the definition of 1-Wasserstein distance. Then, according to the triangle inequality of Wasserstein distance Peyré et al. (2019), we can obtain:
$d_{W}(\mathcal{X}_{(T|f)},\mathcal{X}_{(S^{*}|f)})\leq d_{W}(\mathcal{X}_{(T|f)},\mathcal{X}_{(\hat{X}|f)})+d_{W}(\mathcal{X}_{(\hat{X}|f)},\mathcal{X}_{(S^{*}|f)})$ (11)
Combining Eq. equation 9 and Eq. equation 11 we finished the proof and obtained the Theorem 2.4. By reducing the distance term $d_{W}\left(\mathcal{X}_{(T|f)},\mathcal{X}_{(\hat{X}|f)}\right)$, we have $\mathcal{X}_{(T|f)}\rightarrow\mathcal{X}_{(S^{*}|f)}$. As a result, the expected distance
$\mathbb{E}_{(S^{*}\sim\mathcal{X}_{(S^{*}|f)},T\sim\mathcal{X}_{(T|f)})}\min_{\pi\in\Pi(T,S^{*})}\mathbb{E}_{(x_{T},x_{S^{*}})\sim\pi}d(x_{T},x_{S^{*}})\rightarrow 0,$
with randomly sampling $S^{*}$ and $T$ with $K$ elements.
Appendix N Experiment on Text to Image generation Model (Stable Diffusion)
In this section, we conducted experiments using a stable diffusion model Rombach et al. (2022) fine-tuned with LoRA on a Naruto dataset Cervenka (2022).
Identical Attributes Test (C2). To evaluate the effectiveness of GMValuator on a text-to-image dataset, we use the attributes in the images as ground truth and employ the Identical Attributes Test (C2) to assess its performance. The results presented in Table 14 show that the top $k$ contributors share similar attributes with the generated sample, such as gender, hair, and hat, demonstrating the effectiveness of our approach in evaluating data contributions.
Table 14: Performance of Identical Attributes Test (C2) of some attributes including Gender, Hat, and Hair on Naruto.
Attribute (%) Gender Hat Hair
Top K contributors: $k=3$ $k=4$ $k=5$ $k=3$ $k=4$ $k=5$ $k=3$ $k=4$ $k=5$
GMValuator (l2_distance) 94.33 93.50 92.80 86.30 85.50 84.40 88.00 87.75 88.40
GMValuator (LPIPS) 98.67 98.50 98.80 87.00 86.25 85.40 96.67 96.25 96.00
GMValuator (DreamSim) 99.00 98.50 98.80 86.33 86.75 85.80 98.30 97.75 96.00
Efficiency. To evaluate computational efficiency, we report the time taken to identify significant contributors for a single generated sample in Table 15. The results emphasize the effectiveness and efficiency of our approach for text-image models such as Stable Diffusion. The experiments were conducted in an environment equipped with an RTX A6000 (48GB) GPU and a 12-vCPU Intel(R) Xeon(R) Platinum 8255C CPU @ 2.50GHz.
Table 15: Efficiency using Stable Diffusion on Naruto.
Method GMValuator (l2_distance) GMValuator (LPIPS) GMValuator (DreamSim)
Time (s) 1.27888 3.09580 6.28144
Appendix O Efficiency with Increased Data Size
As a training-free and efficient data valuation approach for generative models, GMValuator is also highly scalable to large-scale and complex datasets. To evaluate this, we assess the efficiency of GMValuator on datasets of varying sizes, illustrating the trend in computational efficiency as the data size increases. In Fig.11, we present the running time versus the dataset size for GMValuator variants (dashed lines, with colors representing different similarity metrics) using the same settings as in Sec.N. This result demonstrates that average running time for various variants of GMValuator on a single generated image, with the recalled number set to 200, is dominated by the re-ranking process. As a result, the running time is nearly independent of the dataset size. Noting that the number of recalled images during the re-rank phase is fixed, thus time cost reveal minimal variation in runtime with increasing dataset size, demonstrating that our method maintains consistent computational efficiency even as the data scale grows.
In Fig.11, we present the time cost for one generated image during the recall phase (No-Rerank), with dataset sizes ranging from 100 K to 50000 K. The results show that the running time remains small across different dataset size. Benefiting from the efficiency of our recall phase, the time overhead increases only marginally. In contrast, as noted in Kong & Chaudhuri (2021), “the run-time complexity of VAE-TracIn is linear with respect to the number of samples, checkpoints, and network parameters”. Similarly, IF4GAN Terashita et al. (2021) is time-intensive, with complexity increasing alongside model size and the number of training epochs, as outlined in Algorithm 2 of their work.
Figure 10: Running time vs. dataset size for variants of GMValuator.
Figure 11: Running time vs. dataset size for recall phase (No-Rerank) of GMValuator.
Figure 12: Running Time vs. Number of Recall Samples
Appendix P Impact of quantization recall
In this section, we evaluate the effect of quantization recall by empirically compare quantization recall with nearest neighbor (NN) search using varying recalled number $K$ during the recall phase. To be specific, We conducted our experiments on the MNIST dataset using a GAN model with recalled number $K$ increasing from 100 to 700. For consistency, we leverage $l_{2}$-distance in the reranking phase.
Efficiency. GMVALUATOR significantly outperforms nearest neighbor search in speed, taking an average of 0.25 seconds to attribute the most significant contributors for a generated sample, which is more than 50 times faster compared to 13.75 seconds for NN search. Whats more, the time cost for NN search will increase linearly with dataset size, while using PQ recall shows only a slight increased time in larger dataset size as shown in Fig 11. The experiments in this section are run in an environment equipped with an RTX A6000 GPU (48GB) and a 12-vCPU Intel(R) Xeon(R) Platinum 8255C CPU @ 2.50GHz.
Effective. As illustrated in Fig. 12, under different recalled numbers (from 100 to 700), the results of GMValuator are similar to the exhaustive KNN search, demonstrating that our method maintains high performance while ensuring computational efficiency.